OpenStudy (anonymous):

-4x-7y=-27 -2x+4y=24

OpenStudy (amistre64):

I take it we need a system of equations.....

OpenStudy (amistre64):

Lets eliminate a variable just to be mean :) -4x -7y =-27 (-2x+4y = 24)(-2) ----------------- -4x -7y = -27 4x -8y = -48 ------------- 0x -15y = -75 -15y = -75 y = -75/-15 = 15/3 = 5 y=5 plug it into one of the equations and solve for "x": -----------------

OpenStudy (amistre64):

you wanna say something....say it :)

OpenStudy (anonymous):

x=2

OpenStudy (amistre64):

x = -2 i think, but I am relying on my head and its batteries are dying

OpenStudy (amistre64):

-2x+4(5) = 24 -2x +20 = 24 -2x = 24 - 20 -2x = 4 x = -2